OpenStudy (anonymous):

Arrange the steps to solve this system of linear equations in the correct sequence. x + y = -2 2x – 3y = -9

OpenStudy (anonymous):

I need the order of the steps from beginning to end

pooja195 (pooja195):

Do you know how to solve it?

OpenStudy (anonymous):

I have a general idea of how to solve it I'm just not completely sure. I mostly just end up putting the steps in the wrong order

pooja195 (pooja195):

ok show me how you solve it then :) show me what order you would put the steps in.

OpenStudy (anonymous):

well I'm pretty sure you have to combine the equations then distribute

pooja195 (pooja195):

hmm an easier approach would be elimination are you familiar with the method?

OpenStudy (anonymous):

no

pooja195 (pooja195):

whats the opposite of 2?

OpenStudy (anonymous):

-2

pooja195 (pooja195):

-2(x + y = -2) 2x – 3y = -9 Step 1: Distribute the -2 -2(x + y = -2)

OpenStudy (anonymous):

The question gives me a list of steps, and I'm supposed to put them in the correct order, so what I really need is someone to help me figure out the sequence of steps

pooja195 (pooja195):

Is there a way you can paste them on to this post?

OpenStudy (anonymous):

yes wait a moment

OpenStudy (anonymous):

a. subtract 3x+3y=-6 (obtained in step one) from 2x-3y=-9 (given) to solve for x b. substitute the value of x in the first equation (x+y=-2) to get y=1 c. The solution for the systems of equations is (-3,1) d. x=-15 e. The solution for the systems of equations is (-15,13) f. add 3x+3y=-6(obtained in step one) to 2x-3y=-9(given), and solve for x g. x=-3 h. substitute the value of x in the first equation (x+y=-2) to get y=13 i. multiply the first equation by 3: 3(x+y)=3(-2) 3X+3Y=-6

OpenStudy (anonymous):

there is a total of 5 steps

pooja195 (pooja195):

O-o huh? :/ i dont get this...

pooja195 (pooja195):

@jim_thompson5910

OpenStudy (anonymous):

we have to figure out the order of these steps

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):



OpenStudy (anonymous):

I don't know

jimthompson5910 (jim_thompson5910):

do you agree that 3y plus -3y = 0y ?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

so what we have to do is multiply everything in equation 1 by 3 to go from x + y = -2 to 3x + 3y = -6

jimthompson5910 (jim_thompson5910):

new system of equations 3x + 3y = -6 2x – 3y = -9

OpenStudy (anonymous):

so that would be the first step? multiply everything by 3?

jimthompson5910 (jim_thompson5910):

yeah in eq 1

OpenStudy (anonymous):

okay, what's next?

jimthompson5910 (jim_thompson5910):

now you add the equations the 3y and -3y add to 0 and go away

jimthompson5910 (jim_thompson5910):

that leaves just x to solve for

OpenStudy (anonymous):

so how would I solve for x?

jimthompson5910 (jim_thompson5910):

what do you get when you add the equations

OpenStudy (anonymous):

5x

jimthompson5910 (jim_thompson5910):

5x equal to what?

OpenStudy (anonymous):

-15

jimthompson5910 (jim_thompson5910):

5x = -15 ----> x = ???

OpenStudy (anonymous):

-3

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

you plug that x value into any equation with x & y in it, then solve for y

OpenStudy (anonymous):

so after adding the two equations what is the next step? I posted the options above

jimthompson5910 (jim_thompson5910):

give it a shot and tell me what the order would be (based off what we've done so far)

OpenStudy (anonymous):

The question doesn't ask for me to solve it. It just wants me to put the steps in order

jimthompson5910 (jim_thompson5910):

you actually do have to solve it

OpenStudy (anonymous):

based on what we've done I'd say the next step would be x=-3

OpenStudy (anonymous):

and then substitute the value of x in the first equation to get y=1

jimthompson5910 (jim_thompson5910):

yes so the solution is (-3,1)

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

np